# Quasiparticle cooling algorithms for quantum many-body state preparation

### v1

@@ II.4.2 Modulated cooling protocol (MCP)
Alternatively, the coupling strength $f_{\tau}$ can be modulated in time such that $F_{h,T}(\epsilon)$ leads to approximately exponential suppression of heating transitions. One particular choice of $f_{\tau}$ is $f_{\tau}=\frac{1}{A\beta}\frac{\sin[\pi(\tau-\tau_{0})/2]}{\sinh[\pi(\tau-\tau_{0})/\beta]},$ (29) where $\tau_{0}=T/2$, $A$ is the normalization factor  [^3] We note that the normalization factor is chosen according to the condition $\sum_{\tau=1}^{T}f_{\tau}=1$. The modulated coupling in Eq. (29) is sign-changing, however, $\sum_{\tau=1}^{T}|f_{\tau}|$ is of order 1 for all cases considered below.. The auxiliary quasienergy is fixed at $\pi h=\pi/2$. We call this approach the modulated cooling protocol (MCP).
In the SCP, transitions occur at the natural resonant frequency of the auxiliary qubits; in the MCP, the resonant frequencies are engineered according to modulation in the time-domain. More specifically, the MCP is designed to satisfy the thermal detailed balance condition, and $F_{h,T}$ takes the form of a broadened step-function (Fermi distribution) in frequency space for $1\ll\beta\ll T$, with ‘thermal’ broadening $\sim 1/\beta$:
$F^{\rm MCP}_{h,T}\approx\pi\bigg{[}\frac{\tanh(\epsilon\beta/2)-\tanh((\epsilon-\pi)\beta/2)}{\tanh(\pi\beta/4)}\bigg{]}.$ (30)
We refer the reader to Appendix A for the derivation of this formula in the limit where $T\to\infty$, and $0<\beta/T\ll 1$, and Appendix B for a more detailed discussion of the detailed balance condition. The frequency-step form of the MCP leads to a suppression
$\bigg{|}\frac{F^{\rm MCP}_{h,T}(+\epsilon)}{F^{\rm MCP}_{h,T}(-\epsilon)}\bigg{|}\approx e^{\epsilon\beta},$ (31)
and heating transitions are now exponentially suppressed. We have also verified that the step-like form of $F^{\rm MCP}_{h,T}(\epsilon)$ is realized for moderate values of $T\approx 10$, which is the practical regime in the presence of noise.
In the rest of this paper, we focus on cooling gapped quasiparticle systems, with $\Delta>0$. As in the case of thermal equilibrium, we anticipate the difficuly of cooling gapless systems to their ground state, since zero energy modes may be excited by arbitrarily small heating elements for any finite $\beta$ (in the case of the MCP). Indeed, on general grounds of the energy-time uncertainty relation, we expect modes of energy $\epsilon$ require times $T\approx\epsilon^{-1}$ to cool.
We note that other, more general protocols can be constructed, which in certain cases may improve cooling efficiency. In particular, $h$ can be modulated during the cooling cycle, as in the adiabatic demagnetization protocol recently studied in Ref. [43]. In Sec. V, when considering cooling of non-integrable systems, we will also vary the coupling strength $\theta$ over cooling cycles in order to reach the steady state more rapidly. Our intention with introducing the two protocols above is to provide a basis for effective cooling, and in particular to highlight the role of modulating the coupling pulse to block unwanted heating.
@@ III Theory of quasiparticle cooling in a spin chain
In this Section, we study the performance of the cooling algorithms using the example of the TFIM, specified by the Floquet operator in Eq. (5). We assume the absence of noise (which we include in Section IV). We first consider the edge auxiliary setup (Fig. 1c). We will see that this leads to solvable dynamics in the sense that the density matrix of the system remains Gaussian. We show that the predictions of the rate equation derived above are in excellent agreement with the exact numerical solution. Further, we study the bulk auxiliary setup (Fig. 1d), where integrability is broken by the coupling to auxiliaries. In this case we compare the rate equation predictions to matrix-product states simulations of the dissipative evolution, finding again a good agreement. We investigate the dominant cooling and heating processes in the two phases, paramagnetic (PM) and antiferromagnetic (AFM), of the TFIM, finding qualitative differences due to the different nature of quasiparticles. By comparing the performance of the SCP and MCP, we find that the latter allows ground state preparation with a fidelity arbitrarily close to 1, in the noiseless case, by increasing the parameter $\beta$.
@@ III.1 Transverse-field Ising model
where $k\in[-\pi;\pi]$ is a quasimomentum. As a function of parameters $J$ and $g$, the model has four phases, separated by phase transition lines at $J=g$, $J=1-g$, where the quasiparticle gap $\Delta$ closes. The four phases are distinguished by the 00 and $\pi$ Majorana edge modes [56, 57, 58]. Here we will be interested in the part of the phase diagram where $20<J,g<1/$ 2, and two phases exist: a paramagnetic (PM) phase at $g>J$, and an antiferromagnetic (AFM) phase at $g<J$. For $g,J\to 0$ these are the phases of the Hamiltonian TFIM; by analogy to this case, we define the Floquet vacuum projector
@@
to project onto states where the valence quasiparticle band is filled, and the conduction band is empty i.e. $\hat{c}_{k}|\Omega_{g}\rangle=0$, $\forall k$. The index $g$ runs over the possible ground states: in the PM phase the vacuum state is unique, while in the AFM phase in the limit $N_{S}\to\infty$, the vacuum state is doubly degenerate, $g=\pm 1$, and the $\mathbb{Z}_{2}$ symmetry is spontaneously broken. In the fermionic counterpart of the TFIM, this degeneracy stems from the existence of zero-(quasi)energy Majorana edge modes, and the two vacuum states of the AFM differ by occupation of the Majorana level. The quasienergy gap between the ground and first excited state is given by $\Delta=2\pi|J-g|$. In the rest of the paper, unless otherwise stated, results for the AFM (PM) phase correspond to the parameter points $J=0.2\ (0.1),\ g=0.1\ (0.2)$.
For the open-boundary chain we consider, the allowed quasimomenta are quantized, $k_{m}\approx\pi(m-1)/N_{S}$, $m=1,2,..,N_{S}$, and the fermionic eigenmodes, defined by creation/annihilation operators $\hat{\eta}_{k}^{\dagger},\hat{\eta}_{k}$, are superpositions of plane waves with wave-vectors ±$\pm k$ (standing waves). The relation between on-site fermionic operators and eigenmodes is
@@ III.2 A solvable model: cooling at the edge
Refer to captionRefer to captionRefer to caption
Figure 2: Performance of edge cooling protocol, for the parameter choice described in the main text. (a) Evolution of quasiparticle density $n(t)$ as a function of rescaled timescale $t=\nu\theta^{2}$, where $\nu$ is total number of cooling cycles, for the PM phase. We compare exact numerical results with $\theta=0.02$ (solid lines) and the rate equation (dashed lines), for both SCP and MCP. (b) Same as in (a) but for AFM phase; (c) Log-fidelity per qubit of the steady state with quasiparticle vacuum, obtain from kinetic equation, as a function of the MCP parameter $T=3\beta$. Black squares are exact numerical data, with $\theta=0.001$, showing that kinetic theory accurately captures the steady-state properties.
We now study a solvable example of cooling in the edge auxiliary setup of Fig 1c. In this case, the auxiliary coupling operator $\hat{\sigma}_{1}^{\pm}$ is free of the Jordan-Wigner string in the fermionic language, being proportional to $c_{1}^{\dagger},c_{1}$ (see Eq. (32)). Therefore, in the rate equation (23), only processes where occupation of one of the quasiparticle levels changes by ±$\pm 1$ are allowed. Due to reflection symmetry, the auxiliary at the other edge contributes equally to the cooling process, and after some algebra, Eq. (23) takes the form:
@@
We will compare the rate equation predictions to exact numerical simulations, which can be performed efficiently in the case of edge cooling: due to the free-fermion nature of the TFIM with boundary auxiliaries, Eq. (1) defines a Gaussian channel, so starting from an initial Gaussian state, such as the identity state, Wick’s theorem guarantees that the complete many-body state information is captured by the $2N\times 2N$-dimensional matrix of 2-point fermion correlators, where $N=N_{S}+2$ is the number of system qubits plus the two auxiliaries. By numerical evolution of the correlation matrix we can assess the accuracy of the rate equation in capturing the two protocols. Further detail on the numerical steps is provided in Appendix E.
As a first characterisation of the cooling performance, we study the dynamics of the quasiparticle density $n(t)=N_{S}^{-1}\sum_{k}n_{k}(t)$. In Fig. 2a,b we plot $n(t)$ for both protocols, SCP and MCP, in the two phases, starting from the maximally mixed state $n(0)=1/2$. We take $T=4$, $h=0.3$ in the SCP, and $T=28=3\beta$, $h=0.5$ in the MCP; the smaller reset time in the SCP case is approximately optimal (see Fig. 2c) while the MCP choice corresponds to $\beta\approx 2\pi\Delta^{-1}$. In both cases $\theta=0.02$ and $N_{S}=20$. We observe a very good agreement between the rate equation theory (dashed lines) and the exact numerical simulation (solid lines). In the considered time window, for both PM and AFM phases under the MCP, the density decays predominantly exponentially in time, with a rate proportional to $\theta^{2}$. In accordance with the arguments regarding heating rates in Section II, the MCP exhibits a saturation of $n(t)$ at a much smaller value than the SCP (though in that case, the saturation is still to a relatively small value on account of the large quasiparticle gap). We also note the long time-scales associated with cooling to states of low quasiparticle density, with $\nu\approx 300\theta^{-2}$ for the PM phase under the MCP and even greater for the AFM. This is a consequence of the edge cooling setup, and the scaling of the Bogoliubov coefficients near the band edge discussed above.
@@
The second line holds in the limit of sufficiently low total quasiparticle density. In this regime, the log-fidelity per qubit and quasiparticle density are approximately synonymous. In Fig. 2c we plot the steady state fidelity for the two pulses as the parameter $\beta$ is increased, with $T=3\beta$ fixed (for the SCP this is just a variation of parameter $T$). For parameters considered, the SCP allows to reach fidelity of order $0.99$ per qubit  while the MCP fidelity (per qubit) is better than one part in $10^{6}$ by the point $\beta\Delta\approx 4\pi$. From Eq. (31), we see that the density of quasiparticles is bounded approximately as $n\sim e^{-\beta\Delta}$. The apparent saturation of MCP around $\beta\Delta=4\pi$ is due to small ringing errors, which can be removed by increasing the ratio $T/\beta$ — we refer the reader to Appendix A for more details. In practice, dominant heating processes arise due to noise (see Sec. IV) and it pays to keep the ratio $T/\beta\gtrsim O(1)$.
@@ III.3 Finite density of auxiliaries
Refer to captionRefer to captionRefer to caption
Figure 3: Performance of the bulk cooling protocol, with one auxiliary per system qubit and $N_{S}=20$. The parameter choice is described in the main text. (a) Quasiparticle density $n(t)$ vs. rescaled timescale $t=\nu\theta^{2}$ in the PM phase. The MPS results (solid lines) are compared to the kinetic theory predictions (dashed lines), for both SCP and MCP. Parameter $\theta=0.1$. (b) Same as in (a) but for the AFM phase. We show MPS results for two coupling strengths, $\theta=0.1$ and $0.05$, with the larger showing corrections to the rate equation. The AFM cools more slowly than the PM, as expected due to the topological nature of the quasiparticles. (c) Log-fidelity per qubit of the steady state with quasiparticle vacuum, calculated from the kinetic equation, as a function of parameter $\beta$, related to the number of unitary layers in one cooling cycle, via $T=5\beta$.
While providing an instructive example, the edge cooling setup is not practical for large systems. We argued above that the cooling timescale diverges with the system size in that case, which holds generally for models with local boundary dissipation [59]. Furthermore, the cooling rate remains constant while the heating rate due to noise scales with system size (see the following Section). Therefore, we now consider the setup in Fig. 1d, with a finite density of auxiliaries. For simplicity, we will assume that there is one auxiliary per system qubit.
In this setup, the dynamics is no longer integrable, and the system’s density matrix develops non-Gaussian correlations due to the fact that the matrix elements for bulk $\sigma_{j}^{\pm}$ operators carry Jordan-Wigner strings in the fermionic formulation (see Eq. (32)). This allows for transitions between many-body eigenstates where the occupation numbers of several quasiparticle levels change at once.
@@
$\sigma^{\pm}_{j;k_{1},\ldots,k_{m};k_{1},\ldots,k_{n}}=\bra{k_{1},\ldots,k_{m}}\hat{\sigma}^{\pm}_{j}\ket{k_{1},\ldots,k_{n}}.$ (45)
@@
Figure 4: Quasiparticle occupation numbers $n^{\infty}_{k}$ in the steady state, as a function of quasienergy $\epsilon_{k}$. We compare MPS simulations (scatter points) and kinetic equation results (red line), for a system with $N_{S}=N_{A}=20$ qubits in (a) PM and (b) AFM phases. The reset time is $T=12=2\beta$, and the coupling $\theta=0.05$. The MPS simulations use a maximum bond dimension of $d=100$. Inset: deviation $|\delta n_{k}|$ between rate equation and MPS values shown in main figure.
@@
In the spin language, these selection rules follow from intuitive arguments given in Ref. [43]. The quasiparticles of the PM phase are magnons, local excitations that can be removed via local spin flips (caused by auxiliary qubit interactions); in contrast, in the AFM phase, quasiparticles are domain-wall-like and can only be removed/created locally in pairs. In Appendix D, we study the distinct behavior of matrix elements in the two phases in detail. In particular, the two-particle matrix elements in the AFM phase are homogeneous in the bulk, but single-particle matrix elements decay exponentially with the distance of the auxiliary qubit from the boundary (since single domain walls can be removed at the edges of the chain), with a decay length set by the correlation length of the model. We will see below that this difference in quasiparticle nature leads to different efficiencies of the cooling in the two phases.
@@
As illustrated in Fig. 3a,b, the theory continues to be accurate for cooling with a finite density of auxiliaries, for both the SCP and MCP protocols. We fix system size $N_{S}=20$ with one auxiliary per site. We use cooling parameters $T=5$, $h=0.3$ in the SCP and $T=12=2\beta$ in the MCP. The coupling is chosen to be $\theta=0.1$ in the PM phase. In the AFM phase, this relatively large coupling leads to visible deviations between the MPS results and the kinetic equation prediction at late times; hence we include simulation results for a weaker coupling, $\theta=0.05$, which leads to a good agreement. We use bond dimension $d=100$ for both the SCP and MCP ($\theta=0.05$) data, and $d=150$ for the MCP ($\theta=0.1$) data. These choices are converged in bond dimension (see Appendix E). Our choice of larger $\theta$ and smaller MCP reset time (compared to the edge cooling case) helps to avoid high entanglement at intermediate times, which affects the accuracy of the MPS simulations. The MCP outperforms the SCP, and in Fig. 3c we provide further confirmation that the MCP can reach much higher fidelities, plotting the increase in the log-fidelity per qubit as the parameter $T=5\beta$ is increased (we choose a large value of $T/\beta$ to remove the ringing artifacts for the range of fidelities shown, see Appendix A).
@@
As a final comparison between the kinetic equation and the MPS numerics, we look at the momentum-resolved quasiparticle occupations in the steady state, in Fig. 4, focusing on the MCP. To reach the steady state in our MPS simulations, we perform the simulations starting from the exact ground state. As we expect the (unique) steady state to be close to the true ground state, this allows for faster real-time simulations compared to starting from a high energy-density state. The figure illustrates that occupation numbers of all levels are below $10^{-2}$, and the remaining population is mostly due to quasiparticles at the band edge. The finite weight on low energy modes stems from the relatively small value of the parameter $\beta\approx 1.2\pi\Delta^{-1}$ chosen to aid the convergence of the MPS simulations.

### v2

@@ II.5.2 Modulated cooling protocol (MCP)
The coupling strength $f_{\tau}$ is modulated in time according to $f_{\tau}=\frac{2}{A\beta}\frac{\sin[\pi(\tau-\tau_{0})/2]}{\sinh[2\pi(\tau-\tau_{0})/\beta]},$ (38) where $\tau_{0}=T/2$, $A$ is the normalization factor  [^3] We note that the normalization factor is chosen according to the condition $\sum_{\tau=1}^{T}f_{\tau}=1$. The modulated coupling in Eq. (38) is sign-changing, however, $\sum_{\tau=1}^{T}|f_{\tau}|$ is of order 1 for all cases considered below., and the auxiliary quasienergy is fixed at $\pi h=\pi/2$. We call this approach the modulated cooling protocol (MCP) .
For finite $T$ and $1\ll\beta\ll T$, the filter function associated to the MCP gives a good approximation to Eq. (37), $F^{\rm MCP}_{h,T}(\epsilon)\approx F_{\beta}(\epsilon)$, with the Fourier series being effectively truncated at order $T$. Hence, the MCP satisfies detailed balance and leads to exponential suppression of heating processes in the limit given above. We have verified that $F^{\rm MCP}_{h,T}(\epsilon)$ in fact closely approximates $F^{\beta}(\epsilon)$ even when $\beta/T\approx 1$ (see Fig. 8, Appendix A). Using the experimental values of Ref. [38], and setting $\gamma T/\theta^{2}\approx 1$, where $\gamma$ is the noise rate and $\theta$ the coupling strength, we find that the current practical regime in the presence of noise is around $T\approx 10$. Reduction of the noise levels in future quantum devices will allow the reset time to be further increased, which as we will see leads directly to higher state preparation fidelities.
In the rest of this paper, we focus on cooling gapped quasiparticle systems, where the many-body gap is $\Delta>0$. Then for high-fidelity ground state preparation, it is sufficient to ensure $\beta\Delta\gg 1$ for the MCP. Due to the detailed balance condition, it is also possible to use the MCP to prepare finite-temperature thermal states, as we review in Appendix B, simply by choosing $\beta$ accordingly. We note that the strategy given above, of first writing down the filter function in the frequency domain and Fourier transforming to obtain the coupling modulation, is general and can be used to target other steady states which are not necessarily in thermal equilibrium.
@@ III Theory of quasiparticle cooling in a spin chain
In this Section, we study the performance of the cooling algorithms using the example of the TFIM, specified by the Floquet operator in Eq. (5). We assume the absence of noise (which we include in Section V). We first consider the edge auxiliary setup (Fig. 1c). We will see that this leads to solvable dynamics in the sense that the density matrix of the system remains Gaussian. We show that the predictions of the rate equation derived above are in excellent agreement with the exact numerical solution. Then, we study the bulk auxiliary setup (Fig. 1d), where integrability is broken by the coupling to auxiliaries. In this case we compare the rate equation predictions to matrix-product states simulations of the dissipative evolution, finding again a good agreement. We investigate the dominant cooling and heating processes in the two phases, paramagnetic (PM) and antiferromagnetic (AFM), of the TFIM, finding qualitative differences due to the different nature of quasiparticles. By comparing the performance of the SCP and MCP, we find that the latter allows ground state preparation with high fidelity, in the noiseless case, by increasing the parameter $\beta$ ($T$).
@@ III.1 Transverse-field Ising model
where $k\in[-\pi;\pi]$ is a quasimomentum. As a function of parameters $J$ and $g$, the model has four phases, separated by phase transition lines at $J=g$, $J=1-g$, where the quasiparticle gap $\Delta$ closes. The four phases are distinguished by the 00 and $\pi$ Majorana edge modes [60, 61, 62]. Here we will be interested in the part of the phase diagram where $20<(J,g)<1/$ 2, and two phases exist: a paramagnetic (PM) phase at $g>J$, and an antiferromagnetic (AFM) phase at $g<J$. For $g,J\to 0$ these are the phases of the Hamiltonian TFIM; by analogy to this case, we define the Floquet vacuum projector
@@
to project onto states where the valence quasiparticle band is filled, and the conduction band is empty i.e. $\hat{\eta}_{k}|\Omega_{g}\rangle=0$, $\forall k$, with the quasiparticle annihilation operators $\hat{\eta}_{k}$ specified below. The index $g$ runs over the possible ground states: in the PM phase the vacuum state is unique, while in the AFM phase in the limit $N_{S}\to\infty$, the vacuum state is doubly degenerate, $g=\pm 1$, and the $\mathbb{Z}_{2}$ symmetry is spontaneously broken. In the fermionic counterpart of the TFIM, this degeneracy stems from the existence of zero-(quasi)energy Majorana edge modes, and the two vacuum states of the AFM differ by occupation of the Majorana level. The quasienergy gap between the ground and first excited state is given by $\Delta=2\pi|J-g|$. In the rest of the paper, unless otherwise stated, results for the AFM (PM) phase correspond to the parameter points $J=0.2\ (0.1),\ g=0.1\ (0.2)$.
For the open-boundary chain we consider, the allowed quasimomenta are quantized, $k_{m}\approx\pi(m-1)/N_{S}$, $m=1,2,..,N_{S}$, and the fermionic eigenmodes, defined by creation (annihilation) operators $\hat{\eta}_{k}^{\dagger}$ ($\hat{\eta}_{k}$), are superpositions of plane waves with wave-vectors ±$\pm k$ (standing waves). The relation between on-site fermionic operators and eigenmodes is
@@
In the paramagnetic limit $J=0$, the quasiparticle annihilation operators are given by $\hat{\eta}_{k}=\frac{1}{\sqrt{N_{S}}}\sum_{j}e^{ijk}\hat{\sigma}_{j}^{+}$. Following the discussion in Subsection II.4 regarding the optimal choice of system-auxiliary coupling in Eq. (8), for the rest of this section we set
where $\mathcal{Q}_{j}$ ($\mathcal{A}_{j}$) represent system (auxiliary) position associated with $j$-th coupling.
@@ III.2 A solvable model: cooling at the edge
We now study a solvable example of cooling in the edge auxiliary setup of Fig 1c. In this case, the system cooling operator $\hat{\sigma}_{1}^{+}$ ($\hat{\sigma}_{1}^{-}$) is free of the Jordan-Wigner string in the fermionic language, being proportional to $\hat{c}_{1}$ ($\hat{c}_{1}^{\dagger}$) (see Eq. (39)). Therefore, in the rate equation (25), only processes where occupation of one of the quasiparticle levels changes by ±$\pm 1$ are allowed. Due to reflection symmetry, the auxiliary at the other edge contributes equally to the cooling process, and Eq. (25) takes the form:
@@
As a first characterisation of the cooling performance, we study the dynamics of the quasiparticle density $n(t)=N_{S}^{-1}\sum_{k}n_{k}(t)$. We compare the rate equation predictions to exact numerical simulations, which can be performed efficiently in the case of edge cooling due to the problem’s free-fermion nature (see Appendix E). In Fig. 2a,b we plot $n(t)$ for both protocols, SCP and MCP, in the two phases, starting from the maximally mixed state $n(0)=1/2$. We take $T=4$, $h=0.3$ in the SCP, and $T=3\beta/2=28$, $h=0.5$ in the MCP; the smaller reset time in the SCP case is approximately optimal (see Fig. 2c) while the MCP choice corresponds to $\Delta\beta\approx 4\pi$. In both cases $\theta=0.02$ and $N_{S}=20$. We observe a very good agreement between the rate equation theory (dashed lines) and the exact numerical simulation (solid lines). In the considered time window, for both PM and AFM phases under the MCP, the density decays predominantly exponentially in time, with a rate proportional to $\theta^{2}$. In accordance with the arguments regarding heating rates in Section II, the MCP exhibits a saturation of $n(t)$ at a much smaller value than the SCP (though in that case, the saturation is still to a relatively small value on account of the large quasiparticle gap). We also note the long time-scales associated with cooling to states of low quasiparticle density, with $\nu\approx 300\theta^{-2}$ for the PM phase under the MCP and even greater for the AFM. This is a consequence of the edge cooling setup, and the scaling of the Bogoliubov coefficients near the band edge discussed above.
@@
The second line holds in the limit of sufficiently low total quasiparticle density. In this regime, the log-fidelity per qubit and quasiparticle density are approximately synonymous. In Fig. 2c we plot the steady state fidelity for the two pulses as the parameter $\beta$ is increased, with $T=3\beta/2$ fixed (for the SCP this is just a variation of parameter $T$). For parameters considered, the SCP allows to reach fidelity of order $0.99$ per qubit  while the MCP fidelity per qubit is better than one part in $10^{6}$ by the point $\beta\Delta\approx 8\pi$. From Eq. (36), we see that the density of quasiparticles is bounded approximately as $n\sim e^{-\beta\Delta}$ for large $\beta$. The apparent saturation of MCP around $\beta\Delta=8\pi$ is due to small ringing errors, which can be removed by increasing the ratio $T/\beta$ — we refer the reader to Appendix A for more details. In practice, dominant heating processes arise due to noise and it pays to keep the ratio $T/\beta\gtrsim O(1)$.
@@
Figure 2: Performance of edge cooling protocol, for the parameter choice described in the main text. (a) Evolution of quasiparticle density $n(t)$ as a function of rescaled timescale $t=\nu\theta^{2}$, where $\nu$ is total number of cooling cycles, for the PM phase. We compare exact numerical results with $\theta=0.02$ (solid lines) and the rate equation (dashed lines), for both SCP and MCP; (b) Same as in (a) but for AFM phase; (c) Log-fidelity per qubit of the steady state with quasiparticle vacuum, obtained from kinetic equation, as a function of the MCP parameter $T=3\beta/2$. Black squares are exact numerical data, with $\theta=0.001$, showing that kinetic theory accurately captures the steady state properties.
@@ III.3 Finite density of auxiliaries
Figure 3: Performance of the bulk cooling protocol, with one auxiliary per system qubit and $N_{S}=20$. The parameter choice is described in the main text. (a) Quasiparticle density $n(t)$ vs. rescaled timescale $t=\nu\theta^{2}$ in the PM phase. The MPS results (solid lines) are compared to the kinetic theory predictions (dashed lines), for both SCP and MCP. Parameter $\theta=0.1$. (b) Same as in (a) but for the AFM phase. We show MPS results for two coupling strengths, $\theta=0.1$ and $0.05$, with the larger showing corrections to the rate equation. The AFM cools more slowly than the PM, as expected due to the topological nature of the quasiparticles. (c) Log-fidelity per qubit of the steady state with quasiparticle vacuum, calculated from the kinetic equation, as a function of the MCP parameter $T=5\beta/2$.
While providing an instructive example, the edge cooling setup is not practical for large systems. We argued above that the cooling timescale diverges with the system size in that case, which holds generally for models with local boundary dissipation [63]. Furthermore, the cooling rate remains constant while the heating rate due to noise scales with system size. Therefore, we now consider the setup in Fig. 1d, with a finite density of auxiliaries. For simplicity, we will assume that there is one auxiliary per system qubit.
In this setup, the dynamics is no longer integrable, and the system’s density matrix develops non-Gaussian correlations due to the fact that the matrix elements for bulk $\hat{\sigma}_{j}^{\pm}$ operators carry Jordan-Wigner strings in the fermionic formulation (see Eq. (39)). This allows for transitions between many-body eigenstates where the occupation numbers of several quasiparticle levels change at once.
@@
$\sigma^{\pm}_{j;k_{1},\ldots,k_{m};k^{\prime}_{1},\ldots,k^{\prime}_{n}}=\bra{k_{1},\ldots,k_{m}}\hat{\sigma}^{\pm}_{j}\ket{k^{\prime}_{1},\ldots,k^{\prime}_{n}}.$ (53)
@@
Figure 4: Quasiparticle occupation numbers $n^{\infty}_{k}$ in the steady state, as a function of quasienergy $\epsilon_{k}$. We compare MPS simulations (scatter points) and kinetic equation results (red line), for a system with $N_{S}=N_{A}=20$ qubits in (a) PM and (b) AFM phases. The reset time is $T=\beta=12$, and the coupling $\theta=0.05$. The MPS simulations use a maximum bond dimension of $d=100$. Inset: deviation $|\delta n_{k}|$ between rate equation and MPS values shown in main figure.
@@
In the spin language, these selection rules follow from the phenomenological discussion in Subsection II.4: since we chose the system-auxiliary couplings in Eq. (43) to match the local magnon excitations of the paramagnetic fixed point, we expect that also within the PM phase single excitations can be removed locally; in contrast, in the AFM phase, quasiparticles are dressed domain walls and can only be removed/created locally in pairs. This intuitive argument was pointed out in Ref. [43]. In Appendix D, we study the distinct behavior of matrix elements in the two phases in detail. In particular, the two-particle matrix elements in the AFM phase are homogeneous in the bulk, but single-particle matrix elements decay exponentially with the distance of the auxiliary qubit from the boundary (since single domain walls can be removed at the edges of the chain), with a decay length set by the correlation length of the model. We will see below that this difference in quasiparticle nature leads to different efficiencies of the cooling in the two phases.
@@
As illustrated in Fig. 3a,b, the theory continues to be accurate for cooling with a finite density of auxiliaries, for both the SCP and MCP protocols. We fix system size $N_{S}=20$ with one auxiliary per site. We use cooling parameters $T=5$, $h=0.3$ in the SCP and $T=\beta=12$ in the MCP. The coupling is chosen to be $\theta=0.1$ in the PM phase. In the AFM phase, this relatively large coupling leads to visible deviations between the MPS results and the kinetic equation prediction at late times; hence we include simulation results for a weaker coupling, $\theta=0.05$, which leads to a good agreement. We use bond dimension $d=100$ for both the SCP and MCP ($\theta=0.05$) data, and $d=150$ for the MCP ($\theta=0.1$) data. These choices are converged in bond dimension (see Appendix E). Our choice of larger $\theta$ and smaller MCP reset time (compared to the edge cooling case) helps to avoid high entanglement at intermediate times, which affects the accuracy of the MPS simulations. The MCP outperforms the SCP, and in Fig. 3c we provide further confirmation that the MCP can reach much higher fidelities, plotting the increase in the log-fidelity per qubit as the parameter $T=5\beta/2$ is increased (we choose a large value of $T/\beta$ to remove the ringing artifacts for the range of fidelities shown, see Appendix A).
@@
As a final comparison between the kinetic equation and the MPS numerics, we look at the momentum-resolved quasiparticle occupations in the steady state, in Fig. 4, focusing on the MCP. To reach the steady state in our MPS simulations, we perform the simulations starting from the exact ground state. As we expect the (unique) steady state to be close to the true ground state, this allows for faster real-time simulations compared to starting from a high energy-density state. The figure illustrates that occupation numbers of all levels are below $10^{-2}$, and the remaining population is mostly due to quasiparticles at the band edge. The finite weight on low energy modes stems from the relatively small value of the parameter $\beta\Delta\approx 2.4\pi$ chosen to aid the convergence of the MPS simulations.